OpenStudy (anonymous):

Convert the following metric measurements. 1576 mL = how many L 352 cm = how many m 26.8 mW = how many W

OpenStudy (anonymous):

1.576 L 3.52 m

OpenStudy (anonymous):

1L = 1000mL So, 1mL = 1/1000 L 1576mL = 1576/1000 L = 1.576L

OpenStudy (anonymous):

1m = 100cm and 1cm = 1/100 m 1W = 1000mW and 1mW = 1/1000 W Use these and do the rest